OpenStudy (linn99123):

The area of the base of a prism is 55 mm2. The perimeter of the base is 30 mm. The height of the prism is 7 mm. What is the surface area of the prism? A. 445 mm2 B. 320 mm2 C. 265 mm2 D. 92 mm2

OpenStudy (ranga):

A prism has identical top and bottom surfaces and it has lateral surface (side surface). Total surface area = 2 * area of base + perimeter * height

OpenStudy (anonymous):

265

OpenStudy (ranga):

Bottom area = 55 Top area will be the same as the bottom. Therefore, 55 Lateral or side surface area = perimeter of base * height = 30 * 7 = 210 Total surface area = 210 + 55 + 55

OpenStudy (anonymous):

320

OpenStudy (anonymous):

320

OpenStudy (anonymous):

Eh?